40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1024     Entry time: Fri Oct 3 15:57:05 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: last night, again 
Last night was basically a repeat of the night before--marginally better locking with the DRMI resonating the +f2
sideband. Several stable locks were achieved, and several control handoffs to DDM signals worked, but never from
lock to lock--that is, a given DD handoff strategy would only work once. This really needs to work smoothly before
more progress can be made.

Also, a 24Hz mode got rung up in one/several of the suspensions--this can also impede the stability of locks.
ELOG V3.1.3-